OpenStudy (anonymous):

how will i get 654 with -81 and 8??

jhonyy9 (jhonyy9):

so there is minus 81 ?

OpenStudy (anonymous):

\[-81\times(8-8-8) +8 - \frac{8}{8}-\frac{8}{8}\]

jhonyy9 (jhonyy9):

why is there (8-8-8) when this is equal (-8)

OpenStudy (anonymous):

Because we don't have -8.

jhonyy9 (jhonyy9):

nice